OpenStudy (anonymous):

find the domain of f(x)=[sqrt(x-4)]/ln(3-2x)+1

OpenStudy (helder_edwin):

u have \[ \large f(x)=\frac{\sqrt{x-4}}{\ln(3-2x)}+1 \]

OpenStudy (helder_edwin):

u have three restrictions: (i) \(x-4\geq0\) (ii) \(3-2x>0\) (iii) \(\ln(3-2x)=0\) solve them separatedly and then intersect.